# Overlap-aware End-to-End Supervised Hierarchical Graph Clustering for Speaker Diarization

### v1

@@ VI-A Comparison with the baseline systems
Table II reports the performance based on the cluster purity and coverage on Voxconverse dev set using pyannote-metric [40]. Cluster purity is defined as the percentage of segments from predicted speakers belonging to the corresponding speaker in the ground truth. Cluster coverage is defined as the percentage of segments from the ground truth speaker covered by the predicted speaker. From the table, it can be observed that the baseline with AHC has high purity but low coverage. However, the baseline with spectral clustering (SC) has lower purity but higher coverage. In our proposed approach, both the purity and the coverage are high, indicating that supervised clustering can achieve the best of both worlds.
Table III shows the false alarm (FA), miss, and confusion error along with the overall DER metric of baseline AHC, SC approaches, and the proposed SHARC/E-SHARC models with and without overlap assignment for three different datasets: AMI Single Distant Microphone (SDM), Voxconverse and DISPLACE challenge datasets. From Table III, it can be observed that SHARC and E-SHARC perform significantly better than AHC and SC baselines on all the datasets. E-SHARC achieves $11\%$ and $13\%$ relative improvements over the best baseline on AMI and Voxconverse datasets, respectively.
@@
From Table III, it can be observed that adding overlap assignment to AHC and SC improves the DER performance. Further improvements are obtained with the proposed SHARC-Overlap and E-SHARC-Overlap approaches. We achieve $11.6\%$, $16.8\%$, and $25.3\%$ relative improvements for AMI, Voxconverse, and DISPLACE, respectively, over the best baseline system.
Table IV shows the results of baseline and proposed approaches with VBx resegmentation [41]. The VBx is initialized with the corresponding clustering output followed by a resegmentation step using the HMM-PLDA model. It can be observed that VBx segmentation improves performance as it helps refine the speaker boundaries based on temporal information. However, SHARC/E-SHARC initialization requires only 1-2 VBx iterations for convergence to the lowest DER. After the VBx iterations, the second speaker is predicted for the overlapping regions based on the second-highest posterior probability. The E-SHARC-Overlap + VBx achieves the lowest DER for all the datasets, resulting in $9.4\%$, $4.3\%$, and $12.6\%$ relative improvements for AMI, Voxconverse, and DISPLACE, respectively over the best baseline system.
TABLE V: DER (%, with overlap $+$ without collar) and DER* (%, without overlap $+$ with 0.25s collar) comparison with state-of-the-art on AMI SDM Eval, Voxconverse Eval, and DISPLACE Eval(phase 2) datasets.
AMI SDM System DER DER*
Pyannote [42] 29.1 -
x-vec+AHC+VBx [41] 27.4 12.6
SelfSup-PLDA-PIC +VBx [9] 23.8 5.5
Raj et al. [43] 23.7 -
Plaquet et al. [44] 22.9 -
GAE-based+ SC [45] - 5.5
GADEC-based [45] - 4.2
E-SHARC (proposed) 19.83 2.9
E-SHARC-Ovp +VBx (prop.) 17.2 2.6
Voxconverse System
Pyannote [42] 11.9 -
Plaquet et al. [44] 10.4 -
GAE-based+ SC [45] - 8.0
GADEC-based [45] - 7.6
E-SHARC (prop.) 11.68 7.6
E-SHARC-Ovp +VBx (proposed) $\bm{10.1}$ 6.3
DISPLACE System
DISPLACE Baseline [36, 46] 32.2 14.6
E-SHARC-Ovp +VBx (prop.) + Baseline SAD 31.4 13.0
Winning system [46] 27.8 7.3
@@
VI-B Comparison with the other published works
Table V compares the proposed approach with state-of-the-art approaches for AMI, Voxconverse, and DISPLACE datasets. Pyannote [42] is the end-to-end pyannote model with overlap detection modules. Raj et al. [43], and Plaquet et al. [44] explore the recent end-to-end models for diarization. The work reported in [45] proposed Graph Attention-Based Deep Embedded Clustering (GADEC), which performs graph attention-based clustering using multi-objective training. It also shows the results of the Graph attentional encoder (GAE) based approach for metric learning followed by spectral clustering. Our E-SHARC-Ovp + VBx outperforms the state-of-the-art results for AMI SDM and Voxconverse systems. In the case of DISPLACE, the winning system investigated different SAD and model combination strategies, while the proposed system was trained on out-of-set (Voxceleb) data.
Figure 4: Plot comparing mean absolute error (MAE) for speaker counting task with $k$ ranging from $20$-$100$ and $\tau\in\{0.0,0.4,0.8\}$ for Voxconverse dev.
Figure 5: 2D t-SNE plot to compare x-vectors and GNN embeddings for two different recordings from Voxconverse dev set. The first column shows the x-vectors with SC labels in different colors, and the second column shows GNN features with ESHARC labels in different colors. Ground truth labels are represented as different shapes. In both cases, the proposed E-SHARC yields representations with better separability. However, the DER deteriorated due to early stopping in Recording 2.
@@ VII-A Choice of hyper-parameters
SHARC parameters ($k$, $\tau$): A high value of $k$ while training allows to capture more edges in the graph and improves the model predictions. However, a higher value of $k$ while testing results in fewer levels and clusters. Therefore, the choice of $k$ is also dependent on the average number of speakers present in a recording. The parameter $\tau$ (edge prediction probability threshold) sets the minimum required probability to allow the edge connection between nodes. Therefore, a higher value leads to fewer connections, generating a higher number of clusters.
Figure 3 shows the impact of different values of $k$ and $\tau$ on DER on the Voxconverse dev set. As $k$ increases, the DER increases as higher $k$ generates fewer speakers. On the other hand, the DER is higher for a smaller value of $\tau$ ($\tau<0.2$). The best values of $k$ and $\tau$ are $30$ and $0.8$.
@@ VII-B Speaker counting task
We also perform the speaker counting task evaluation using the proposed model. Figure 4 shows the performance of the speaker counting task for the Voxconverse dataset. It shows the impact of $k$ and $\tau$ on the mean absolute error (MAE) between the ground truth number of speakers and the predicted number of speakers. The figure is divided into three categories of recordings based on the number of speakers present. Plot (a) shows the MAE for recordings with a higher number of speakers ($>10$). As $k$ increases, MAE increases, and as $\tau$ increases, MAE reduces. Plot (b) shows the MAE for recordings containing a moderate number of speakers ($5-10$). Finally, plot (c) shows the MAE for recordings containing a small number of speakers ($1-4$). In this case, MAE is relatively low across different $k$ and $\tau$.
@@
Figure 6: 2D scatter plot showing no. of ground truth speakers vs no. of predicted speakers using (a) SC and (b) ESHARC on Voxconverse dev set. The different colors represent different ranges of average DER (%) for a pair of (#ground speaker, #predicted speakers).
VII-C Representation visualization
@@
Figure 5 shows the t-SNE plots for two recordings (two rows) from Voxconverse dev set for baseline and the proposed ESHARC approach. The first column shows a t-SNE plot of x-vectors with SC, and the second column shows GNN features with ESHARC. Recording 1 contains five speakers represented by different shapes. The SC predicts only three speakers out of five, as shown in different colors. However, E-SHARC can form five different clusters and provide a lower DER. It can also be observed that the within-cluster covariance is lower for the GNN features compared to x-vectors. For recording 2, the SC can predict the correct number of speakers. On the other hand, E-SHARC splits spk1 from ground truth into two different speakers shown in different colors. Although the features are well separated in terms of speaker clusters, the stopping criterion has resulted in early stopping.
@@
Figure 6 shows a 2D scatter plot comparing the performance of SC and ESHARC in terms of the predicted number of speakers on Voxconverse dev. The color represents different ranges of filewise DER for the particular pair of (#groundtruth speakers, #predicted speakers). It can be observed that, for files with a lower number of speakers ($<4$), the model predicts a very high number of speakers, which leads to very high DERs ($>50$) for SC (Figure 6 (a)). On the other side, Figure 6 (b) shows the scatter plot for the E-SHARC model. The E-SHARC model can correctly predict speakers for files with fewer speakers in the ground truth, which results in lower DER (0-10). The E-SHARC predicts fewer speakers than the ground truth for other files, but it does not affect the DER performance adversely.
@@ VIII Conclusion
This paper proposes a novel end-to-end supervised hierarchical neural graph clustering approach for speaker diarization. The proposed approach involves supervised graph model training with a clustering-based loss. This model, called E-SHARC, performs hierarchical clustering using a GNN model by minimizing intra-speaker distances and maximizing inter-speaker distances. The model is further extended to perform overlapped speaker prediction. This enables the prediction and assignment of multiple speakers in the overlapped speech regions, further improving the diarization performance. We also showed the effectiveness of the proposed approach on a challenging code-mixed DISPLACE dataset.
@@ Acknowledgments
The authors would like to thank Amrit Kaul for helping in the initial experiments. The authors would like to thank Michael Free, Rohit Singh, and Shakti Srivastava from BT Research for their valuable input. The authors also thank Hervé Bredin for providing pretrained pyannote SAD and Overlap detection models.

### v2

@@ VI-A Comparison with baseline systems
Table II reports the performance based on the cluster purity and coverage on Voxconverse dev set using pyannote-metric [51]. Cluster purity is defined as the percentage of segments from predicted speakers belonging to the corresponding speaker in the ground truth. Cluster coverage is defined as the percentage of segments from the ground truth speaker covered by the predicted speaker. From the table, it can be observed that the baseline with AHC has high purity but low coverage. However, the baseline with spectral clustering (SC) has lower purity but higher coverage. In our proposed approach, both the purity and the coverage are high, indicating that supervised clustering can achieve better trade-off.
Table III shows the false alarm (FA), miss, and confusion error along with the overall DER metric of baseline AHC, SC approaches, and the proposed SHARC/E-SHARC models with and without overlap assignment for three different datasets: AMI Single Distant Microphone (SDM), Voxconverse and DISPLACE challenge datasets. From Table III, it can be observed that SHARC and E-SHARC perform significantly better than AHC and SC baselines on all the datasets. E-SHARC achieves $11\%$, $13\%$ and $20\%$ relative improvements over the best baseline on AMI, Voxconverse and DISPLACE datasets, respectively.
@@
Figure 3: 2D t-SNE plot to compare x-vectors and GNN embeddings for two different recordings from Voxconverse dev set. Ground truth labels are represented as shapes while predictions are represented as colors. In both cases, the proposed E-SHARC yields representations with improved separability. However, the DER for recording-2 deteriorated due to early stopping.
From Table III, it can be observed that adding overlap assignment improves the DER performance for all models. We achieve $11.6\%$, $5.2\%$ , and $19.8\%$ relative improvements for AMI, Voxconverse, and DISPLACE, respectively, over the best baseline system.
Table IV shows the results of baseline and proposed approaches with VBx resegmentation [43]. The VBx is initialized with the corresponding clustering output of the SHARC/ESHARC models. This is followed by a re-segmentation step using the HMM-PLDA model. It can be observed that VBx segmentation improves performance as it helps refine the speaker boundaries based on temporal information. After the VBx iterations, the second speaker is predicted for the overlapping regions based on the second-highest posterior probability. The E-SHARC-Overlap + VBx achieves the lowest DER for all the datasets, resulting in $9.4\%$, $4.3\%$, and $12.6\%$ relative improvements for AMI, Voxconverse, and DISPLACE, respectively over the best baseline.
@@
VI-B Comparison with other published works
Table V compares the proposed approach with state-of-the-art approaches for AMI, Voxconverse, and DISPLACE datasets using different VAD configurations. Pyannote [28] is the end-to-end pyannote model with overlap detection modules. Raj et al. [45], and Plaquet et al. [46] explore the recent end-to-end models for diarization. The work reported in [44] proposed Graph Attention-Based Deep Embedded Clustering (GADEC), which performs graph attention-based clustering using multi-objective training. It also shows the results of the Graph attentional encoder (GAE) based approach for metric learning followed by spectral clustering. Our E-SHARC-Ovp + VBx outperforms the state-of-the-art results for AMI SDM and Voxconverse systems using oracle VAD, although this setting is somewhat deprecated in the recent evaluations. We have also reported performance of systems of VoxSRC 2023 challenge [47, 48] for Voxconverse datasets, which involves more complex models and larger training data for diarization. We have reported results of ResNet x-vector models with system VAD, VBx and OSD [20]. We have also compared with recent Perceiver based end-to-end diarization system (DiaPer) [20], trained with $2500$ hours of simulated data as opposed to $75$/$400$ hours of real/simulated data for AMI/Voxconverse in the proposed approach. In the case of DISPLACE, the winning system investigated different VAD and model combination strategies, while the proposed system was trained on out-of-set (Voxceleb) data. The results suggest that the E-SHARC approach performs better than the DiaPER EEND system and x-vec + VBx system in case of AMI SDM system. This may be attributed to the improved modeling of the intra and inter speaker relationships through the graph clustering. However, Plaquet et al. achieves the best results in this setting. In that work, the authors use a combination of EEND and clustering jointly. In such a case, the EEND performs efficient modeling within a small window with overlap handling and VAD, while a linking/clustering across windows is performed by a spectral clustering setting. As a future scope, it is worthwhile to explore the incorporation of EEND with our proposal of graph clustering. We hypothesize that such a framework could further improve the results and combine the advantages of graphs with EEND setting.
Figure 5: Comparison of mean absolute error (MAE) for speaker counting task for different values of $k$ and $\tau$ on Voxconverse Dev. data
@@ VII-A Choice of hyper-parameters
SHARC parameters ($k$, $\tau$): A high value of $k$ while training allows to capture more edges in the graph and improves the model predictions. The parameter $\tau$ (edge prediction probability threshold) sets the minimum required probability to allow the edge connection between nodes. Figure 4 shows the impact of different values of $k$ and $\tau$ on DER on the Voxconverse dev set. As $k$ increases, the DER increases as higher $k$ generates fewer speakers. On the other hand, the DER is higher for a smaller value of $\tau$ ($\tau<0.2$). The best values of $k$ and $\tau$ are $30$ and $0.8$.
@@ VII-B Speaker counting task
We also perform the speaker counting task evaluation using the proposed model. Figure 5 shows the performance of the speaker counting task for the Voxconverse dataset. It shows the impact of $k$ and $\tau$ on the mean absolute error (MAE) between the ground truth number of speakers and the predicted number of speakers. The figure is divided into three categories of recordings based on the number of speakers present. Plot (a) shows the MAE for recordings with a higher number of speakers ($>10$). As $k$ increases, MAE increases, and as $\tau$ increases, MAE reduces. Plot (b) and (c) shows the MAE for recordings containing a small number of speakers where the trends are different. In these settings, MAE is relatively low across different $k$ and $\tau$. A higher value of $k$ while testing results in fewer levels and clusters, while a higher value of $\tau$ leads to fewer connections and generates a higher number of clusters. Therefore, for plots (a) and (b), lower $k$ and higher $\tau$ leads to better MAE, but for plot (c) the trend is different where the best values of $k$ and $\tau$ are relatively higher and lower, respectively.
@@
Figure 6: 2D scatter plot showing # of ground truth speakers vs # of predicted speakers using (a) AHC, (b) SC, (c) SHARC and (d) E-SHARC models on Voxconverse Dev. data. The different colors represent different ranges of average DER (%).
VII-C GNN vs LSTM architectures
Similar to GNNs, other neural architectures like LSTM may also be capable of performing speaker embedding extraction and clustering [52]. We have experimented with bidirectional LSTM (BLSTM) architecture, where all the other data and modeling settings are kept identical to the proposal of SHARC. We replaced the GNN layer with BLSTM layer in the SHARC model such that number of parameters are similar. The results, shown in Table VI, indicate that the GNN based SHARC performs better than the BLSTM architecture for the AMI diarization task. This experiment illustrates the improved representation learning achieved using the GNN for speaker diarization.
TABLE VI: Performance comparison of GNN vs BLSTM architecture in SHARC on AMI SDM Eval set using oracle VAD.
Architecture # Hid. Dim # Parameters DER %
GNN 2048 8M 21.27
BLSTM 2048 16M 22.92
BLSTM 1024 8M 22.13
VII-D Representation visualization
@@
Figure 3 shows the t-SNE plots for two recordings (two rows) from Voxconverse dev set for baseline and the proposed E-SHARC approach. The first column shows a t-SNE plot of x-vectors with SC, and the second column shows GNN features with E-SHARC. Recording 1 contains five speakers represented by different shapes. The SC predicts only three speakers out of five, as shown in different colors. However, E-SHARC can form five different clusters and provide a lower DER. It can also be observed that the within-cluster covariance is lower for the GNN features compared to x-vectors. For recording 2, the SC can predict the correct number of speakers. On the other hand, E-SHARC splits spk1 from ground truth into two different speakers shown in different colors. Although the features are well separated in terms of speaker clusters, the stopping criterion has resulted in early stopping.
@@
Figure 6 shows a 2D scatter plot comparing the performance of AHC, SC, SHARC and E-SHARC in terms of the predicted v/s true number of speakers on Voxconverse dataset. It can be observed that, the AHC and the SC predict more number of speakers compared to ground-truth resulting in high DERs ($>50$%). On the other side, the SHARC and the E-SHARC models are more accurate in predicting the true number of speakers, resulting in a lower file-wise DERs.
@@ VIII Conclusion
This paper extends our prior work [14] and provides a detailed analysis of end-to-end supervised hierarchical neural graph clustering approach for speaker diarization. The proposed approach involves supervised graph model training with a clustering-based loss. This model, called E-SHARC, performs hierarchical clustering using a GNN model by minimizing intra-speaker distances and maximizing inter-speaker distances. The model is further extended to perform overlapped speaker prediction. This enables the prediction and assignment of multiple speakers in the overlapped speech regions, further improving the diarization performance. We also showed the effectiveness of the proposed approach on a challenging code-mixed DISPLACE dataset.
@@ Acknowledgments
The authors would like to thank Amrit Kaul for helping in the initial experiments. The authors would like to thank Michael Free, Rohit Singh, and Shakti Srivastava from BT Research for their valuable input. The authors also thank Hervé Bredin for providing pretrained pyannote VAD and Overlap detection models.